OpenStudy (anonymous):

Does anyone know how to solve a system of equations with 3 variables?

OpenStudy (saiyan):

what is the equation

OpenStudy (saiyan):

if you have windows and your on an online school you can use snipping tool to take a pic of it and post it here

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

x+2y-z=-3 2x-2y+2z=8 2x-y+3z=9

OpenStudy (saiyan):

you can use substitution or elimination

OpenStudy (amistre64):

if we define x in terms of y and z, we can use that in the other 2 equations ... and it reduces to a system of 2 equations

OpenStudy (anonymous):

Ok @amistre64

OpenStudy (anonymous):

so now what do I do? @rishavraj

rishavraj (rishavraj):

ok see lets use elimination method.....we will eliminate "x" x + 2y - z = -3 --> say 1st equation 2x - 2y + 2z = 8 which is also x - y + z = 4 ----> say 2nd equation 2x - y + 3z = 9 --> say 3rd equation

OpenStudy (anonymous):

Ok, I'm getting it:)

rishavraj (rishavraj):

now multiply 2nd equation from -1 and add the new equation with 1st equation..... "x" gets eliminated...... did u understand??

OpenStudy (anonymous):

Yes! :D

OpenStudy (anonymous):

Thanks:)

rishavraj (rishavraj):

so whts the equation u got??

OpenStudy (anonymous):

I got 3x+z=5 and 5x+5z=15 for the 2nd equation

OpenStudy (anonymous):

Was I wrong?

rishavraj (rishavraj):

hmmmmmmm -_- wht u did ???? i askd to eliminate "x" the 1st equation is x + 2y - z = -3 and when u multiply 2nd equation from -1 u get -x + y - z = -4now add these two equations

OpenStudy (anonymous):

3y=-7

OpenStudy (anonymous):

Oh! I used subsitution!

rishavraj (rishavraj):

its 3y - 2z = -7 --->say 4th equation

OpenStudy (anonymous):

Ok:)

rishavraj (rishavraj):

hey hold on...... wht were the values u got...the answer ??

rishavraj (rishavraj):

bco if u meant x = 1 y = -1 and z = 2 then thts correct :)) @SydneyMay17

OpenStudy (anonymous):

Yes! I did! <3 Thanks

OpenStudy (anonymous):

@rishavraj Sorry, I had to eat dinner

rishavraj (rishavraj):

oo cool... not a big deal :))

OpenStudy (anonymous):

Can I ask another question, if I open a new link?

rishavraj (rishavraj):

for sure

OpenStudy (anonymous):

kk!